# A.X K1 Technical Report

### v1

@@ 3.2 Pre-training Process
    Long-Context Adaptation: To enable long-context capabilities, we adopt a multi-stage curriculum that progressively extends the model’s context length from 4K to 16K and then to 32K, using a total of 600 billion tokens. This is motivated by the demands of agentic tasks, which require long-horizon reasoning, planning, and action across extended contexts. Accordingly, the long-context stages include data with long chains of thought and repository-level code. We apply learning rate decay during the 16K extension phase and keep it constant during the final 32K training stage.
@@ 5.2 Evaluation Results of Thinking Mode
Domain Benchmark Lang. A.X K1 (519B-A33B) DeepSeek-V3.1 (685B-A37B) GLM-4.6 (357B-A32B) Knowledge KMMLU Korean 80.2 76.5 79.9 KMMLU-Redux Korean 77.9 75.9 78.2 KMMLU-Pro Korean 68.1 71.4 71.8 CLIcK Korean 84.9 84.5 84.9 KoBALT Korean 48.8 59.7 59.2 MMLU-Pro English 81.5 85.1 82.9 GPQA Diamond English 74.0 77.9 78.0 Instruction Following IFBench (prompt-loose) English 64.7 41.5 43.4 IFEval (prompt-strict) English 80.4 84.4 86.1 IFEval-ko (prompt-strict) Korean 81.0 79.2 85.8 Math AIME25 English 89.8 88.4 86.0 AIME25-ko Korean 80.4 81.3 80.4 HRM8K Korean 84.3 84.3 83.9 Code LiveCodeBench v6 (25.2.1 - 25.5.1) English 75.8 69.5 76.0 LiveCodeBench-ko Korean 73.1 66.2 55.9 HumanEval+ English 87.2 86.0 83.5 HumanEval+ ko Korean 90.2 93.9 86.0 MBPP+ English 93.0 99.2 98.9 SciCode English 32.4 39.1 38.4 Long Context AA-LCR English 36.0 53.3 54.3 Humanity's Last Exam English 7.2 13.0 13.3 Agent $\tau^{2}$ Telecom English 58.1 37.4 70.5

### v3

@@ 3.2 Pre-training Process
    Long-Context Adaptation: To enable long-context capabilities, we adopt a multi-stage curriculum that progressively extends the model’s context length from 4K to 16K and subsequently to 32K, using a total of 600 billion tokens. This design is motivated by the demands of agentic tasks, which require long-horizon reasoning, planning, and action over extended contexts. Accordingly, the long-context training stages incorporate data featuring long chains of thought and repository-level code. We maintain a constant learning rate throughout all extension phases.
@@ 5.2 Evaluation Results of Thinking Mode
Domain Benchmark Lang. A.X K1 (519B-A33B) DeepSeek-V3.1 (685B-A37B) GLM-4.6 (357B-A32B) Knowledge KMMLU Korean 80.2 76.5 79.9 KMMLU-Redux Korean 77.9 75.9 78.2 KMMLU-Pro Korean 68.1 71.4 71.8 CLIcK Korean 84.9 84.5 84.9 KoBALT Korean 48.8 59.7 59.2 MMLU-Pro English 81.5 85.1 82.9 GPQA Diamond English 74.0 77.9 78.0 Instruction Following IFBench (prompt-loose) English 64.7 41.5 43.4 IFEval (prompt-strict) English 80.4 84.4 86.1 IFEval-ko (prompt-strict) Korean 81.0 79.2 85.8 Math AIME25 English 89.8 88.4 86.0 AIME25-ko Korean 80.4 81.3 80.4 HRM8K Korean 84.3 84.3 83.9 Code LiveCodeBench v6 (25.2.1 - 25.5.1) English 75.8 69.5 76.0 LiveCodeBench-ko Korean 73.1 66.2 55.9 HumanEval+ English 87.2 86.0 83.5 HumanEval+ ko Korean 90.2 93.9 86.0 MBPP+ English 93.0 99.2 98.9 SciCode English 32.4 39.1 38.4 Long Context AA-LCR English 36.0 53.3 54.3 Humanity's Last Exam English 8.6 13.0 13.3 Agent $\tau^{2}$ Telecom English 58.1 37.4 70.5